OpenStudy (anonymous):

maximize A=xy if x+y=12

OpenStudy (anonymous):

y=16-2x so put that within the equation A=x(16-2x) A=16x-2x^2 taking the derivative of a you get A'(x)=16-4x when A'= 0 you get -16=-4x 4=x if x = 4 y=16-2(4)= 8 x=4 y=8